A power station has an output of 800 MW and the thermal efficiency is 28 per cent. Determine the coal consumed in tonne per hour if the calorific value of the coal is 31 MJ/kg
1
Expert's answer
2020-12-14T07:32:30-0500

Answer "#"

Energy consumption is given by

"Input=efficiency\times \times output"

Input=0".28\times80MW"

"Q=22.4MJ"

Coal consumption is

"m=\frac{Q}{sT}=\frac{22.4MJ}{273\times31MJ/Kg}=2.7Kg"